# Revisiting the Canonicalization for Fast and Accurate Crystal Tensor Property Prediction

### v3

@@ A.1 Related work
GNN-Based Methods. CGCNN is a pioneering GNN model specifically designed for handling crystal structures [47]. This model proposed to represent crystal structures as multi-edge crystal graphs. It was applied to predict various single-value properties such as formation energy and band gap. Since then, several GNN methods have been developed to improve upon CGCNN through exploring various network designs or leveraging prior knowledge [1, 26, 2, 3, 21]. These GNN methods are primarily designed for single-value property prediction and do not address the prediction of high-order tensor properties, such as dielectric or elastic tensors. Furthermore, they lack the ability to preserve the equivariance required for accurate high-order tensor property prediction. In contrast, recent studies attempted to ensure equivariance through specialized network architectures [28, 25, 11, 46, 34, 51, 53]. These approaches generally employ harmonic decomposition to achieve equivariance for tensor properties. Within these network architectures, many operations are required, such as tensor products and combining irreducible representations. These processes significantly increase computational costs, especially when handling higher-order data.
@@
Transformer-Based Methods. Transformers, with their self-attention mechanism and parallel processing capabilities, are particularly well-suited for predicting crystal material properties. Matformer [49], one of the earliest Transformer-based networks used for crystal material property prediction, encoded crystal periodic patterns by using the geometric distances between the same atoms in neighboring unit cells. This addressed the issue in earlier GNN-based methods, including CGCNN, MEGNet, GATGNN, and others, which neglected the periodic patterns of infinite crystal structures. Subsequently, more advanced Transformer-based approaches were proposed. Some methods, such as ComFormer [50], CrystalFormer [44], and CrystalFormer [42], typically incorporate either enhanced graph construction techniques or physical priors, exhibiting impressive results in the prediction of single-value properties. Furthermore, DOSTransformer [20] is tailored for the density of states prediction, utilizing prompt-guided multi-modal transformer architecture to achieve super performance. Nevertheless, these models are not directly applicable for accurate high-order tensor property prediction due to their inability to capture the necessary equivariance.
@@
Equivariant Learning. Recently, two general-purpose equivariant learning techniques for the arbitrary group have gained widespread attention: frame averaging [22, 36, 5] and canonicalization [27, 17, 6]. Frame averaging is developed based on group averaging and involves averaging over smaller frames instead of the entire group. Canonicalization, on the other hand, is inherently related to frame averaging and can be viewed as a more general form of it. [27]. In particular, Lin et al. [22] introduced minimal frame averaging, a method designed to achieve both efficient and exact equivalence in equivariant learning. This approach is primarily applicable to scenarios where the input and output spaces are identical, such as the n-body problem. It does not extend to cases involving different spaces, such as crystal tensor prediction. If this method were extended to tensor property prediction, it would represent a special case of Proposition 3.4, where both orbit canonicalization and rigid registration are achieved using QR decomposition. Within Proposition 3.4, however, different functions can be employed for orbit canonicalization and rigid registration. For instance, the Kabsch algorithm [18, 19] can be employed for rigid registration, while polar decomposition can be utilized for canonicalization.
@@ A.2 Details of Framework Architecture
Crystal Graph Construction. In GoeCTP, we use the crystal graph construction from Yan et al. [50, 51] to describe the structure and relationships within crystals. Specifically, we first convert the fractional coordinate system $(\mathbf{A},\mathbf{F},\mathbf{H})$ into the Cartesian coordinate system $(\mathbf{A},\mathbf{X},\mathbf{H})$ as introduced in Section 2. Then, we assume that the output crystal graph is represented as $\mathcal{G}(\mathcal{V},\mathcal{E})$, $\mathcal{V}$ denotes the set of nodes $v_{i}$ in the crystal graph, where each node $v_{i}$ contains atomic features $\mathbf{v}_{i}=(\bm{a}_{i},\hat{\bm{p}}_{i})$. $\mathcal{E}$ represents the set of edges denoted as $e_{ij}$, which are typically constructed based on the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$. When the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$ is less than a given radius $R$, i.e. $d_{ij}=||\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}||_{2}\leq R$, an edge $e_{ij}$ will be built with edge feature $\mathbf{e}_{ij}=\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}$. Here, $R$ is established based on the distance to the $k$-th nearest neighbor, and different values of $\bm{k}=[k_{1},k_{2},k_{3}]\in\mathbb{Z}^{3}$ represent different edges between nodes $v_{i}$ and $v_{j}$. Since there are no equivariance requirements for subsequent models, any other graph construction methods can be used to replace this part, such as Wang et al. [44], Yan et al. [50].
eComFormer. The eComFormer has demonstrated strong performance across a range of single-value property prediction tasks [50]. It is built on an SO(3)-equivariant crystal graph representation, where the interatomic distance vectors are employed to represent the edge features of the graph. The model converts this crystal graph into embeddings and utilizes a transformer architecture, incorporating both a node-wise transformer layer and a node-wise equivariant updating layer, to extract rich geometric information during the message-passing process.
Figure 3: The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer, adapted from Yan et al. [50].
@@
$\bm{\alpha}_{ij}=\frac{\bm{q}_{ij}\circ\bm{\xi}_{K}(\bm{k}_{ij})}{\sqrt{d_{\bm{q}_{ij}}}},\bm{msg}_{ij}=\operatorname{sigmoid}(\operatorname{BN}(\bm{\alpha}_{ij}))\circ\bm{\xi}_{V}(\bm{\upsilon}_{ij}),$ (8)
@@
where $\bm{\xi}_{K}$, $\bm{\xi}_{V}$ represent nonlinear transformations applied to key and value features, respectively. The operators $\circ$ denote the Hadamard product, $\operatorname{BN}(\cdot)$ refers to the batch normalization layer, and $\sqrt{d_{\bm{q}_{ij}}}$ indicates the dimensionality of $\bm{q}_{ij}$. Then, node feature $\bm{f}_{i}$ is updated as follows,
$\bm{m}s\bm{g}_{i}=\sum_{j\in\mathcal{N}_{i}}\bm{m}s\bm{g}_{ij},\bm{f}_{i}^{\mathrm{new}}=\bm{\xi}_{msg}(\bm{f}_{i}+\mathrm{BN}(\bm{m}s\bm{g}_{i})),$ (9)
where $\bm{\xi}_{msg}(\cdot)$ denoting the softplus activation function.
The node-wise equivariant updating layer is designed to effectively capture geometric features by incorporating node feature $\mathbf{a}_{i}$ and edge feature $||\mathbf{e}_{ji}||_{2}$ as inputs and stacking two tensor product (TP) layers [7]. It uses node feature $\bm{f}_{i}^{l}$ and equivalent vector feature $\mathbf{e}_{ji}$ embedded by corresponding spherical harmonics $\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})=c_{0}$,$\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})=c_{1}*\frac{\mathbf{e}_{ji}}{||\mathbf{e}_{ji}||_{2}}\in\mathbb{R}^{3}$ and $\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{5}$ to represent the input features. Gathering rotational information from neighboring nodes to the central node $i$, the first TP layer is shown as
$\bm{f}_{i,0}^{l}=\bm{f}_{i}^{l^{\prime}}+\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})),\bm{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})),\lambda\in\{1,2\},$ (10)
@@
where $\bm{f}_{i}^{l^{\prime}}$ is linearly transformed from $\bm{f}_{i}^{l}$, $|\mathcal{N}_{i}|$ and $\mathbf{TP}_{\lambda}$ represent the number of neighbors of node $i$ and TP layers with rotation order $\lambda$ respectively. Then, to represent the invariant node features, the second TP layer is written as
$\bm{f}_{i}^{l*}=\frac{1}{|\mathcal{N}_{i}|}(\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,0}^{l},\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,1}^{l},\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,2}^{l},\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji}))),$ (11)
@@
stacking the two tensor product layers together using both linear and nonlinear transformations, the output $\bm{f}_{i,updated}^{l}$ is combined as
$\bm{f}_{i,updated}^{l}=\bm{\sigma}_{\mathrm{equi}}(\mathrm{BN}(\bm{f}_{i}^{l*}))+\mathrm{LN}_{\mathrm{equi}}(\bm{f}_{i}^{l}),$ (12)
@@
with $\bm{\sigma}_{\mathrm{equi}}$ denoting a nonlinear transformation made up of two softplus layers with a linear layer positioned between them. The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer are shown in Fig. 3.
@@ A.3 Proofs
A.4 Proofs of Lemma A.1
Lemma A.1.
@@
According to Definition 3.2, for any $g\in O(3)$ and $\mathbf{M}\in\mathcal{V}$, we have $C_{\mathbf{M}}(\mathbf{M})=\mathbf{M_{0}}$ and $C_{\mathbf{M}}(g\cdot\mathbf{M})=\mathbf{M_{0}}$. Thus,
@@
A.5 Proofs of Lemma A.2
Lemma A.2.
@@
According to Definition 3.3, $\mathbf{M}_{1}=R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})\cdot\mathbf{M}_{2}$. Thus, for any $g\in O(3)$, we have
@@
A.6 Proof of Proposition 3.4
Proposition 3.4. ($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we can define a new function $h(\mathbf{M})=R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$, such that $h(\mathbf{M})$ is $O(3)$ equivariant for tensor prediction.
@@
According to Lemma A.2 and Lemma A.1, for any $g\in O(3)$ we have
@@
A.7 Proof of Proposition 3.6
Proposition 3.6. (Orbit Canonicalization and Rigid Registration for Crystal Tensor Prediction.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$ of a crystal $\mathbf{M}$, we can define functions $f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$ and $f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q}$, where $f_{p1}(\cdot)$ and $f_{p2}(\cdot)$ correspond to performing polar decomposition on the lattice matrix. In this perspective, $f_{p1}(\cdot)$ serve as the orbit canonicalization function, while $f_{p2}(\cdot)$ serve as the rigid registration function. The matrix $\mathbf{H}$ is thus identified as the canonical form.
@@
Given the lattice matrices $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$, where $\mathbf{Q^{\prime}}$ is an arbitrary $O(3)$ group transformation, applying polar decomposition on both $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$ yields the following expressions:
$\mathbf{L}=\mathbf{Q}\mathbf{H}$ (17)
$\mathbf{Q^{\prime}}\mathbf{L}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (18)
By combining Eq. 17 and 18, we can have the following equation:
$\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (19)
Because $\mathbf{Q^{\prime}}\mathbf{Q}(\mathbf{Q^{\prime}}\mathbf{Q})^{\top}=\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{Q}^{\top}\mathbf{Q^{\prime}}^{\top}=\mathbf{I}$, it follows that $\mathbf{Q^{\prime}}\mathbf{Q}$ is an orthogonal matrix. According to the uniqueness of polar decomposition, we can have $\mathbf{Q^{\prime}}\mathbf{Q}=\mathbf{Q_{1}}$, which implies $\mathbf{Q_{1}}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}$. If $\mathbf{H_{1}}\neq\mathbf{H}$, this would violate the uniqueness property of polar decomposition. Therefore, we conclude that $\mathbf{H_{1}}=\mathbf{H}$, indicating that $\mathbf{H}$ remains unaffected by the $O(3)$ group transformation. Consequently, the crystal configuration $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$ can serve as the canonical form. Performing polar decomposition on the lattice matrix to obtain $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$, i.e. $f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$, can thus be regarded as the orbit canonicalization function. Furthermore, based on Eq. 17, performing polar decomposition on the lattice matrix to obtain $\mathbf{Q}$, i.e. $f_{p2}(\mathbf{M},\mathbf{M_{0}})=\mathbf{Q}$, can be regarded as rigid registration function, as shown below:
@@
A.8 $O(3)$ equivariance for crystal tensor properties

### v4

@@ A.1 Related work
GNN-Based Methods. CGCNN is a pioneering GNN model specifically designed for handling crystal structures (xie2018crystal). This model proposed to represent crystal structures as multi-edge crystal graphs. It was applied to predict various scalar properties such as formation energy and band gap. Since then, several GNN methods have been developed to improve upon CGCNN through exploring various network designs or leveraging prior knowledge (chen2019graph; louis2020graph; choudhary2021atomistic; das2023crysgnn; lin2023efficient). These GNN methods are primarily designed for scalar property prediction and do not address the prediction of high-order tensor properties, such as dielectric or elastic tensors. Furthermore, they lack the ability to preserve the equivariance required for accurate high-order tensor property prediction. In contrast, recent studies attempted to ensure equivariance through specialized network architectures (mao2024dielectric; lou2024discovery; heilman2024equivariant; wen2024equivariant; pakornchote2023straintensornet; yanspace; zhong2023general). These approaches generally employ harmonic decomposition to achieve equivariance for tensor properties. Within these network architectures, many operations are required, such as tensor products and combining irreducible representations. These processes significantly increase computational costs, especially when handling higher-order data.
@@
Transformer-Based Methods. Transformers, with their self-attention mechanism and parallel processing capabilities, are particularly well-suited for predicting crystal material properties. Matformer (yan2022periodic), one of the earliest Transformer-based networks used for crystal material property prediction, encoded crystal periodic patterns by using the geometric distances between the same atoms in neighboring unit cells. This addressed the issue in earlier GNN-based methods, including CGCNN, MEGNet, GATGNN, and others, which neglected the periodic patterns of infinite crystal structures. Subsequently, more advanced Transformer-based approaches were proposed. Some methods, such as ComFormer (yancomplete), CrystalFormer (wang2024conformal), CrystalFormer (taniaicrystalformer), and CrystalFramer (ito2025rethinking) typically incorporate either enhanced graph construction techniques or physical priors, exhibiting impressive results in the prediction of scalar properties. Furthermore, DOSTransformer (lee2024density) is tailored for the density of states prediction, utilizing prompt-guided multi-modal transformer architecture to achieve super performance. Nevertheless, these models are not directly applicable for accurate high-order tensor property prediction due to their inability to capture the necessary equivariance.
@@
Equivariant Learning. Equivariance can be achieved through specialized architectural designs, such as incorporating high-degree steerable features (han2025survey). However, due to the structural constraints imposed by equivariant architectures, these models are often difficult to optimize (manolache2025learning; xie2025tale), and their complex designs may limit the universal approximation property (pacini2025universality; cenuniversally). Recently, two general-purpose techniques for achieving equivariant learning under arbitrary groups have gained widespread attention: frame averaging (linequivariance; punyframe; duval2023faenet) and canonicalization (ma2024canonicalization; kaba2023equivariance; dymequivariant). These methods represent promising alternatives to address the limitations of conventional equivariant models, as they impose no architectural constraints. Frame averaging is developed based on group averaging and involves averaging over smaller frames instead of the entire group. Canonicalization, on the other hand, is inherently related to frame averaging and can be viewed as a more general form of it. (ma2024canonicalization). In particular, Lin et al. (linequivariance) introduced minimal frame averaging, a method designed to achieve both efficient and exact equivalence in equivariant learning. This approach is primarily applicable to scenarios where the input and output spaces are identical, such as the n-body problem. It does not extend to cases involving different spaces, such as crystal tensor prediction. If this method were extended to tensor property prediction, it would represent a special case of Proposition 1, where both orbit canonicalization and rigid registration are achieved using QR decomposition. Within Proposition 1. However, different functions can be employed for orbit canonicalization and rigid registration. For instance, the Kabsch algorithm (kabsch1976solution; lawrence2019purely) can be employed for rigid registration, while polar decomposition can be utilized for canonicalization.
@@ A.2 Details of Framework Architecture
[…]
eComFormer. The eComFormer has demonstrated strong performance across a range of scalar property prediction tasks (yancomplete). It is built on an SO(3)-equivariant crystal graph representation, where the interatomic distance vectors are employed to represent the edge features of the graph. The model converts this crystal graph into embeddings and utilizes a transformer architecture, incorporating both a node-wise transformer layer and a node-wise equivariant updating layer, to extract rich geometric information during the message-passing process.
Figure 3: The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer, adapted from yancomplete.
[…]
@@
$\boldsymbol{\alpha}_{ij}=\frac{\boldsymbol{q}_{ij}\circ\boldsymbol{\xi}_{K}(\boldsymbol{k}_{ij})}{\sqrt{d_{\boldsymbol{q}_{ij}}}},\boldsymbol{msg}_{ij}=\operatorname{sigmoid}(\operatorname{BN}(\boldsymbol{\alpha}_{ij}))\circ\boldsymbol{\xi}_{V}(\boldsymbol{\upsilon}_{ij}),$ (7)
@@
where $\boldsymbol{\xi}_{K}$, $\boldsymbol{\xi}_{V}$ represent nonlinear transformations applied to key and value features, respectively. The operators $\circ$ denote the Hadamard product, $\operatorname{BN}(\cdot)$ refers to the batch normalization layer, and $\sqrt{d_{\boldsymbol{q}_{ij}}}$ indicates the dimensionality of $\boldsymbol{q}_{ij}$. Then, node feature $\boldsymbol{f}_{i}$ is updated as follows,
$\boldsymbol{m}s\boldsymbol{g}_{i}=\sum_{j\in\mathcal{N}_{i}}\boldsymbol{m}s\boldsymbol{g}_{ij},\boldsymbol{f}_{i}^{\mathrm{new}}=\boldsymbol{\xi}_{msg}(\boldsymbol{f}_{i}+\mathrm{BN}(\boldsymbol{m}s\boldsymbol{g}_{i})),$ (8)
where $\boldsymbol{\xi}_{msg}(\cdot)$ denoting the softplus activation function.
[…]
$\boldsymbol{f}_{i,0}^{l}=\boldsymbol{f}_{i}^{l^{\prime}}+\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\boldsymbol{f}_{j}^{l^{\prime}},\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})),\boldsymbol{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\boldsymbol{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})),\lambda\in\{1,2\},$ (9)
@@
where $\boldsymbol{f}_{i}^{l^{\prime}}$ is linearly transformed from $\boldsymbol{f}_{i}^{l}$, $|\mathcal{N}_{i}|$ and $\mathbf{TP}_{\lambda}$ represent the number of neighbors of node $i$ and TP layers with rotation order $\lambda$ respectively. Then, to represent the invariant node features, the second TP layer is written as
$\boldsymbol{f}_{i}^{l*}=\frac{1}{|\mathcal{N}_{i}|}(\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\boldsymbol{f}_{j,0}^{l},\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\boldsymbol{f}_{j,1}^{l},\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\boldsymbol{f}_{j,2}^{l},\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji}))),$ (10)
@@
stacking the two tensor product layers together using both linear and nonlinear transformations, the output $\boldsymbol{f}_{i,updated}^{l}$ is combined as
$\boldsymbol{f}_{i,updated}^{l}=\boldsymbol{\sigma}_{\mathrm{equi}}(\mathrm{BN}(\boldsymbol{f}_{i}^{l*}))+\mathrm{LN}_{\mathrm{equi}}(\boldsymbol{f}_{i}^{l}),$ (11)
@@
with $\boldsymbol{\sigma}_{\mathrm{equi}}$ denoting a nonlinear transformation made up of two softplus layers with a linear layer positioned between them. The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer are shown in Fig. 3.
@@ A.3 Proofs
Proofs of Lemma 1
Lemma 1.
@@
According to Definition 2, for any $g\in O(3)$ and $\mathbf{M}\in\mathcal{V}$, we have $C_{\mathbf{M}}(\mathbf{M})=\mathbf{M_{0}}$ and $C_{\mathbf{M}}(g\cdot\mathbf{M})=\mathbf{M_{0}}$. Thus,
@@
Proofs of Lemma 2
Lemma 2.
@@
According to Definition 3, $\mathbf{M}_{1}=R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})\cdot\mathbf{M}_{2}$. Thus, for any $g\in O(3)$, we have
@@
Proof of Proposition 1
Proposition 1. ($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we can define a new function $h(\mathbf{M})=R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$, such that $h(\mathbf{M})$ is $O(3)$ equivariant for tensor prediction.
@@
According to Lemma 2 and Lemma 1, for any $g\in O(3)$ we have
@@
Proof of Proposition 2
Proposition 2. (Orbit Canonicalization and Rigid Registration for Crystal Tensor Prediction.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$ of a crystal $\mathbf{M}$, we can define functions $f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$ and $f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q}$, where $f_{p1}(\cdot)$ and $f_{p2}(\cdot)$ correspond to performing polar decomposition on the lattice matrix. In this perspective, $f_{p1}(\cdot)$ serve as the orbit canonicalization function, while $f_{p2}(\cdot)$ serve as the rigid registration function. The matrix $\mathbf{H}$ is thus identified as the canonical form.
@@
Let $\mathbf{L}\in\mathbb{R}^{3\times 3}$ be an invertible lattice matrix. By the polar decomposition, there exist a unique orthogonal $\mathbf{Q}\in O(3)$ and a unique symmetric positive-definite $\mathbf{H}\in\mathbb{R}^{3\times 3}$ such that
$\mathbf{L}=\mathbf{Q}\mathbf{H},\qquad\big(\mathbf{L}^{\top}\mathbf{L}\big)^{1/2}=\big(\mathbf{H}^{\top}\mathbf{H}\big)^{1/2}=\big(\mathbf{H}\mathbf{H}\big)^{1/2}=\mathbf{H}.$ (16)
Consider an arbitrary $O(3)$ group transformation $\mathbf{Q^{\prime}}\in O(3)$ acting on the lattice, i.e., $\mathbf{L}\mapsto\mathbf{Q^{\prime}}\mathbf{L}$. Apply polar decomposition to $\mathbf{Q^{\prime}}\mathbf{L}$:
$\mathbf{Q^{\prime}}\mathbf{L}=\mathbf{Q}_{1}\mathbf{H}_{1},\qquad\mathbf{H}_{1}=\big((\mathbf{Q^{\prime}}\mathbf{L})^{\top}(\mathbf{Q^{\prime}}\mathbf{L})\big)^{1/2}.$ (17)
Since $(\mathbf{Q^{\prime}})^{\top}\mathbf{Q^{\prime}}=\mathbf{I}$, we have
$\mathbf{H}_{1}=\big(\mathbf{L}^{\top}\mathbf{L}\big)^{1/2}=\mathbf{H}.$ (18)
Therefore, $\mathbf{H}$ remains unaffected by the $O(3)$ group transformation. Because $\mathbf{A}$ and $\mathbf{F}$ are unchanged by $O(3)$ group transformation and $\mathbf{H}$ is invariant by (18), the triplet
$\mathbf{M}_{0}\coloneqq f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$ (19)
is identical for all representatives of the $O(3)$-orbit of $\mathbf{M}$. Therefore, $f_{p1}(\cdot)$ serves as an orbit canonicalization map and the matrix $\mathbf{H}$ is the canonical form.
Furthermore, based on $\mathbf{L}=\mathbf{Q}\mathbf{H}$, performing polar decomposition on the lattice matrix to obtain $\mathbf{Q}$, i.e. $f_{p2}(\mathbf{M},\mathbf{M_{0}})=\mathbf{Q}$, can be regarded as rigid registration function, as shown below:
@@
A.4 $O(3)$ equivariance for crystal tensor properties